OpenStudy (anonymous):

If the distance covered by an object in time t is given by s(t) = 2t2 + 3t, where s(t) is in meters and t is in seconds, what is the average velocity over the interval from 2 seconds to 4 seconds? 15 meters/second 14 meters/second 13 meters/second 12 meters/second 11 meters/second

OpenStudy (pratyush5):

Do you know differential calculus ?

OpenStudy (anonymous):

no im new to this??

OpenStudy (anonymous):

s=2t^2+3t Differentiate the above equation with respect to t..we get, ds/dt=4t+3 We know that ds/dt=velocity Therefore, v=4t+3 Put different values for t and get the answer..all the best :)

OpenStudy (pratyush5):

@sid0830 has done exactly what I said. So basically you don't know diffrentiation ? right ?

OpenStudy (anonymous):

i got a difference of 8?

OpenStudy (pratyush5):

@davidza6496 I need you to confirm me one thing. Did you get how @sid0830 arrived at 4t + 3 ? If not there is another easy way of doing this.

OpenStudy (anonymous):

yeah i understand it now!! thanks

OpenStudy (pratyush5):

So now plug in the values of t = 2 and t = 4. So what values of 'v' do you get ?

OpenStudy (anonymous):

i got 11 for 2 and 19 for 4

OpenStudy (anonymous):

@pratyush5

OpenStudy (pratyush5):

Brilliant now we know average velocity = v1 + v2 whole divide by 2. So we get 11+19/2 = 10/2 = 15